E/D Night UT 08-Mar-2020

Plan for E/D Night UT 08-Mar-2020

(Rollover from UT 03-Mar-2020)
  • Engineering Tasks
    • SOUL SX AO
      • On-sky checkouts of SX
      • OSCO checkouts - at least 2x
      • ESM data collection
        • Files located: OBS-N: /home/lbto/Service_Observing/Eng_D_Mar_2020/ESM_Testing/
          • M67 twilight until 2am
            • R=9.4 and R=15.1 AO Ref Stars
          • M92 2am until twilight
            • R=12.2 and R=16.3 AO Ref Stars
          • Arp 143 twilight until 2am
            • R=14.7 AO Ref Star
          • Hoag's Object midnight until twilight
            • R=13.4 AO Ref Star
    • General Telescope
      • PCS patch 20200226 for TCS 2020B - in place
        • If any problems, roll back patch - no troubleshooting
        • PCS patch 20200226 for TCS 2020B release is available (but not yet activated). This patch hopefully fixes the problem of co-pointing error after instrument authorization changes (see IT#7380). To activate the patch, do the following in a command-line window:

          1. ssh tcs@tcs1.mountain.lbto.org
          2. cd /lbt/tcs/current
          3. ./tcs-build latest PCS (it is OK to answer "n" to the restart question. This way, patch will be run after TO restart affected subsystem.)

          In case of any problem, please revert it back. To revert the patch, enter the following commands:

          1. ssh tcs@tcs1.mountain.lbto.org
          2. cd /lbt/tcs/current
          3. ./tcs-build revert PCS
    • LBC Testing
      • Mixed-mode LBC-R/LUCI-1 tests - dithering tests - medium priority - 0.5 hr - presets worked in full binocular mode - can be done in poor weather
      • Trailing test
      • Photometric standards - requires clear skies but any seeing
    • LUCI Bino or Mono
      • Photometric standards - requires clear skies but any seeing
      • Files located: OBS-N: /home/lbto/Service_Observing/Eng_D_Mar_2020/LUCI_PhotZP/
    • MODS
      • Photometric standards - requires clear skies but any seeing
  • Science Targets
    • Novae Targets - MODS-Bino Dual Grating - 0.75 hours per target (M. Wagner)
      • Files located: OBS-N: /home/lbto/Service_Observing/Eng_D_Mar_2020/Novae_Wagner
        • V392 Per 04:43:21.37 47:21:25.87
          V2860 Ori 06:09:57.453 12:12:25.15
          V959 Mon 06:39:38.61 05:49:26.74
          V435 CMa 07:13:45.84 -21:12:31.3
    • Triple Merger Candidates - LUCIs-mono or bino (B. Rothberg)
      • Files located: OBS-N: /home/lbto/Service_Observing/Eng_D_Mar_2020/Triple_AGNs_Rothberg_LD_2019B_007
      • LUCI observations have priority over MODS
      • Triple Merger Candidate SDSSJ 0838+5055 - visible twilight untl 2am (hits 30deg limit)
        • LUCI-1 - 2.75 hrs G200 1" slit HKspec
        • MODS-Bino - 1 hr Dual Grating 1" slit
        • Requires 1" or better on AGW/WFS
      • Triple Merger Candidate SDSSJ 0853+5245- visible twilight untl 2am (hits 30deg limit)
        • LUCI-1 - 2.75 hrs G200 1" slit HKspec
        • MODS-Bino - 1 hr Dual Grating 1" slit
        • Requires 1" or better on AGW/WFS
      • Triple Merger Candidate SDSS J1342+19133- all 2nd half - higher priority
        • LUCI-1 - 2.75 hrs G200 1" slit HKspec
        • MODS-Bino - 1 hr Dual Grating 1" slit
        • nuclei separations are 5" and 8" - so looser constraints
    • NGC 2403-5 - MODS-Bino -1.5-2.5 hr some clouds seeing < 1.5" (M. Edwards)
      • Files located: OBS-N: /home/lbto/Service_Observing/Eng_D_Mar_2020/Eta_Carina_Edwards/
  • Poor Weather Backup Targets
    • B. Rothberg
      • Local Mergers - MODS Bino or LUCI (mono or bino)
        • Files are located on OBS-N: /home/lbto/Service_Observing/Eng_D_Mar_2020/Rothberg_COBandhead_LD_2019B_005/
        • MODS (dual grating)
          • UGC 4079 - 07:55:06 +55:42:13 - 1 hour
          • NGC 3310 - 10:38:45 +53:30:11 - 1 hour
          • NGC 2744 - 09:04:39 +18:27:52 - 1 hour
          • NGC 6598 - 18:08:56 +69:04:04 - 1 hour (end of night)
        • LUCI (mono or bino)
          • Files are located on OBS-N: /home/lbto/Service_Observing/Eng_D_Mar_2020/Rothberg_COBandhead_LD_2019B_005/
          • Arp 187 - 05:04:53 -10:14:52 - visible only until 9pm (covers zJ and HKspec observations) - G200 HK/zJspec observations
          • Arp 156 - 10:42:38 +77:29:41 (requires the CWL 2.3 cals) G210 K-band - visible most of the night - relative distance to telluric is the real limitation
      • ICRF3 project with USNO (LUCI mono or bino)
        • targets available all night
        • Files Located on OBS-N: /home/lbto/Service_Observing/Eng_D_Mar_2020/ICRF3_Rothberg_2020A/
        • Will need to export LUCI XML files from OT file in directory depending on config used (i.e. L1, L2, or Fraternal Bino)
      • H-band Library (poorest conditions) - LUCI mono or bino
        • Files Located on OBS-N: /home/lbto/Service_Observing/Eng_D_Mar_2020/H-band_Library/
          • LUCI xml scripts need to be exported from OT xml file
        • STAR: RA Dec Class Hmag
        • HD 20797 03 24 40.55 +64 35 09.57 M0Iab 0.5
        • HD 42543 06 12 19.10 +22 54 30.7 M2I 0.9
        • HD 44537 06 24 53.90 +49 17 16.4 M2I 0.8
        • BD -15 1319 06 15 01.88 -15 52 49.7 K5II 6.4
        • HD 169305 18 31 32 +49 07 17 M2III 1.0
  • Closed Dome Tasks
    • LUCI-1 or LUCI-2 checks - 0.5-1hr (D. Thompson) -if needed
    • SX SOUL NCPA data collection
      • RetroReflector (RR) to be installed IF night is called or before 8pm
      • If RR not availble - use ARGOS cal unit - but susceptible to vibrations



Log for E/D Night UT 08-Mar-2020

00:00 Ran up LUCIs and init'd all mechanisms. Then put in N30 + N30_FS on LUCI1. Got a camwheel error in moving from N3.75 to N30. I initialized the camera from the Instrument GUI and tried the move again; this time it worked. Later Dave reminded me that after such a sequence - failed move, init from IMGUI, successful move - the FM4 offsets will have been applied 2x. Indeed that was the case. Realigned the fold mirror ~01:30 UT. L1 0076 is image after adjustment of FM4.

01:15-01:30 Check field alignments. L1 needed adjustment (see above; 0076 is final image). L2 needs still a bit of adjustment. Images 0055 and 0056 show a shadow at top left corner. Will have to investigate in a bit. It is not bad - just the corner.

01:30 Dave called and talked me through PFSalign on L1. Because of the repeated application of the FM offsets from N375 to N30 after the failed move, FM4 was far off and pupil alignment had to be checked. Done now. Final pupil image is 82 and final field image is 81.

05:30 Still closed. Dave clarified the procedure for the AO patrol stage limits (IT 8026) and Shane and I will attempt that task.

05:44 Configuring for measuring AO patrol field limits:
  • Deploy ARGOS cal source: Stow pins in, HBS on. Apply last set of M2 Global Offsets used with ARGOS cal source.
  • L1 images 83, 84 out of focus (ellipitcal donut). N375, H, min DIT, N30_FS and FC off. Checked LUCI1 by init'ing all units and toggling back and forth from H -> K -> H.
  • Closed loop to remove coma & other aberr.
  • Rotator measurements:
    • 87 341 deg
    • 88 251 deg
    • 89 161 deg
    • 90 71 deg
  • Star traces out a very elongated ellipse - even with uncorrected flexure taken into account.
  • Try 2 of rotator measurements with loops closed.
    • set rotator max vel to 1 deg/sec.
    • putting FeII since we nearly saturated with H
    • loops closed
    • 91 341 min exp with FeII does not show up FS edges. Could go longer, but I'll switch to H
    • 92 341 H
    • 93 251 H
    • 94 161 H (Shane closing shutter a bit to decrease intensity of source)
    • 95 71 H
  • Set up a 270 x 270 subwin to take 0.33 sec and take a cube with loops closed and rotator slewing from 0->360 deg.
  • Now loops remain closed but up DIT to 2.51 to see FS edges. Take exposures at 360, 270, 180, 90, 0 with loops closed throughout rotator slews.
    • 106 360 deg
    • 107 270 deg
    • 108 180 deg
    • 109 90 deg
    • 110 0 deg
  • The above set allows for flex correction. The ellipse traced out is similar to what we saw in the 'movie' (105).
  • Now Shane moved the hexapod to offset the ARGOS light source and paused the loop. (Set rot speed back to 5 deg/s)
    • 111 0 deg
    • 112 90 deg
    • 113 180 deg
    • 114 270 deg
    • 115 360 deg
  • The above dataset (111-115) looks like it will give a good rotator center measurement.
  • Next, find the stage position that puts the light source at 1024,1024.
    • Shane moved hexapod back to original position
    • 117 with rotator at park and N375
    • 118 with N30 instead at H
    • 119 at K - slight misalignment now.
    • moved FM4 to align pinholes - off by about 50 pix in x and 30 in y.
    • 123 spot at 1195, 1150
    • Shane is moving the hexapod to try to get the spot on 1024, 1024.
    • 124 1300, 1281 opposite of desired motion.
    • 125 945, 1111
    • 126 900, 1000 --- starting over (moving only the stages now, not the hexapod)
    • 127 792, 1000
    • 128 1013, 1003
    • 129 1068, 975
    • 130 1034, 1033 --- can only move hexapod by 0.1mm and even then, doesn't go exactly to commanded position
    • 131 1030, 1038
    • 133 1024.07 1026.59
    • 134 1023.94 1024.94 --- psf is very elongated - vibration?
    • 135 1023.91 1023 --- stage positions for this are -77.1, -51.98
    • In the previous set, from 126, Shane had the loop closed, and turned down high order gains so he could move the stage while keeping the loops closed, thus moving the secondary in this way.
    • Shutting down occam & resting shell.

-- BarryRothberg - 04 Mar 2020
Topic revision: r9 - 08 Mar 2020, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback